OpenStudy (anonymous):

I WILL MEDAL BEST ANSWER What is the value of x? 4(6 + 7) = 4 • 6 + x • 7

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

\[24+28=4\times6+x \times7\]

OpenStudy (anonymous):

\[24+28=24+x \times7\] \[24-24+28=x \times7\] \[28/7=x\]

OpenStudy (anonymous):

28/7=4

OpenStudy (anonymous):

distribute on the left side... then look at the left side and compare to the right... x should be obvious

OpenStudy (anonymous):

Thank you guys! Would you mind helping me with a few more?